# OccRWKV: Rethinking Efficient 3D Semantic Occupancy Prediction with Linear Complexity

### v1

@@ I INTRODUCTION
Building upon these insights, we introduce OccRWKV, the first RWKV-based 3D semantic occupancy network. In contrast to previous networks that jointly learn semantics and occupancy, OccRWKV separates these predictions into distinct branches. This separation facilitates specialized learning within each domain, enhancing prediction accuracy and fully leveraging the complementary properties of semantic and geometric features in the subsequent feature fusion stage. We integrate novel Sem-RWKV, Geo-RWKV, and BEV-RWKV blocks into these branches to capture long-distance dependencies critical for semantic accuracy and occupancy prediction. Furthermore, by projecting features into the BEV space, we reduce feature fusion latency, enabling real-time inference without compromising performance.
@@
OccRWKV is high-performence. OccRWKV achieves state-of-the-art performance (mIoU = 25.1) on the SemanticKITTI benchmark. (§ IV-B)
@@ II RELATED WORK
3D semantic occupancy prediction is crucial for interpreting occluded environments, as it discerns the spatial layout beyond visual obstructions by merging geometry with semantic clues. The field has seen diverse approaches, broadly categorized into CNN-based and Transformer-based methods. CNN-based methods have demonstrated proficiency in inferring occupancy from various inputs. The Co-Occ [7] framework adopts a multi-modal strategy that fuses LiDAR and camera data, enhanced by volume rendering regularization and a Geometric- and Semantic-aware Fusion module, achieving notable performance on public benchmarks. LowRankOcc [19] employs tensor decomposition and low-rank recovery to address spatial redundancy, leading to state-of-the-art results on multiple datasets. Other notable works such as JS3C-Net [20] and SSC-RS [9] adeptly manage the complexity of outdoor scenes using point cloud data. Transformer-based methods leverage the attention mechanism for feature aggregation and have shown promising results. TPVFormer [6] introduces a tri-perspective approach that combines BEV with two additional planes, achieving LiDAR-like perception using camera inputs alone.
@@ III-A Semantic Segmentation Branch
Voxelization Layer: In this paper, the 3D surrounding scene to be predicted is divided by voxels. The semantic branch consists of a voxelization layer and three encoder Sem-RWKV Blocks with the same structures. The voxelization layer converts an input point cloud , spanning a range of , into sparse voxel features with a dense spatial resolution of . Each point within the point cloud is discretized to a voxel index using the following equation [9]:
@@ III-B Occupancy Prediction Branch
Geo-RWKV Block: The Occupancy Prediction Branch (Fig. 3) commences with an input layer employing a kernel and comprises three Geo-RWKV blocks as the encoder. Each Geo-RWKV block maintains a consistent architecture, integrating a residual block with a VRWKV module and a BEV projection module. Each VRWKV module incorporates a Spatial Mixing module and a Channel Mixing module, adhering to the formulations and operations delineated in the Sem-RWKV block description.
The residual block initially processes the voxels obtained from the point cloud, and its output voxel features serve as the input to the BEV projection module. The dense 3D features are aligned along the -axis, and 2D convolutions are applied to compress the feature dimensions, yielding dense Geometric-BEV features . By leveraging the VRWKV module’s capability to capture long-range dependencies with linear complexity, the occupancy prediction branch can effectively process and refine the geometric information within the voxels to generate enhanced Geometric-BEV features. These enhanced features are subsequently utilized for feature fusion.
@@ III-C BEV Feature Fusion Branch
The BEV feature fusion branch employs a U-Net-like structure with 2D convolutions and BEV-RWKV blocks. Specifically, the encoder (Fig. 4) has an initial input layer and four down-sampling stages, each paired with a BEV-RWKV block. The concatenation of Semantic-BEV features and Geometric-BEV features is first fed into the input layer and then into the first BEV-RWKV block. Before the next BEV-RWKV block, an ARF module from [9] takes the previous stage’s output and semantic-bev and geometric-bev representations at the same scale as inputs and outputs the fused features containing informative semantic context and geometric structure. The decoder uses up-sampling and skip connections to recover spatial details, culminating in a final convolution that outputs the 3D semantic occupancy grid , with representing the number of classes.
@@ V CONCLUSIONS
In conclusion, OccRWKV, our novel network, successfully addresses the challenge of balancing performance and efficiency in 3D semantic occupancy prediction. It delivers state-of-the-art accuracy with a mIoU of 25.1 on the SemanticKITTI benchmark and maintains efficient real-time performance at 22.2 FPS. The network’s scalability makes it a robust solution for practical applications in robot navigation and autonomous driving. Field deployments confirm OccRWKV’s effectiveness in real-world settings, validating its suitability for future integration in complex environments. The video of OccRWKV can be found in the supplementary material.

### v3

@@ I INTRODUCTION
Building upon these insights, we introduce OccRWKV, the first RWKV-based 3D semantic occupancy network. Unlike earlier approaches that combine semantic and occupancy predictions, OccRWKV uses separate pathways for each task. This design allows each branch to focus on its specific learning objectives, improving both semantic and geometric predictions. The system then combines these features effectively in a later fusion step. The architecture includes specialized RWKV blocks for semantics, geometry, and bird’s-eye-view processing, which help capture important relationships across different parts of the input. By converting features to a bird’s-eye-view format, the system can process data efficiently while maintaining high accuracy.
@@
OccRWKV is high-performance. OccRWKV achieves state-of-the-art performance (mIoU = 25.1) on the SemanticKITTI benchmark. (§ IV-B)
@@ II RELATED WORK
3D semantic occupancy prediction [10] is crucial for interpreting occluded environments, as it discerns the spatial layout beyond visual obstructions by merging geometry with semantic clues. The field has seen diverse approaches, broadly categorized into CNN-based and Transformer-based methods. CNN-based methods have demonstrated proficiency in inferring occupancy from various inputs. The Co-Occ [7] framework adopts a multi-modal strategy that fuses LiDAR and camera data, enhanced by volume rendering regularization and a Geometric- and Semantic-aware Fusion module, achieving notable performance on public benchmarks. LowRankOcc [19] employs tensor decomposition and low-rank recovery to address spatial redundancy, leading to state-of-the-art results on multiple datasets. Other notable works such as JS3C-Net [20] and SSC-RS [9] adeptly manage the complexity of outdoor scenes using point cloud data. Transformer-based methods leverage the attention mechanism for feature aggregation and have shown promising results. TPVFormer [6] introduces a tri-perspective approach that combines BEV with two additional planes, achieving LiDAR-like perception using camera inputs alone.
@@ III-A Semantic Segmentation Branch
Voxelization Layer: We partition the 3D environment into voxels for prediction. The semantic component employs a voxelization layer followed by three Sem-RWKV Blocks of identical structure. Our system transforms an input point cloud within the range into voxel features , creating a spatial resolution of . For each point , we calculate its voxel index [9] using:
@@ III-B Occupancy Prediction Branch
Geo-RWKV Block: The occupancy prediction pathway (Fig. 3) is initialized with a convolutional layer, which is succeeded by a cascade of three Geo-RWKV blocks functioning as the encoder. These blocks exhibit a uniform architectural configuration, incorporating a residual connection that amalgamates both VRWKV and BEV projection modules. The VRWKV component executes spatial and channel mixing operations in accordance with the methodology established in the Sem-RWKV framework.
The residual structure commences by processing voxels derived from point cloud inputs, generating voxel representations that subsequently serve as the input for the BEV projection module. The three-dimensional dense features undergo alignment along the -axis, followed by the application of two-dimensional convolutions for dimensional reduction, resulting in a set of dense Geometric-BEV features . Through the exploitation of the VRWKV module’s capacity for efficient long-range dependency modelling with linear computational complexity, the occupancy prediction pathway effectively processes and enhances the geometric information encoded within the voxel representation to generate refined Geometric-BEV features. These enhanced representations are subsequently utilized in the feature fusion process.
@@ III-C BEV Feature Fusion Branch
The BEV feature fusion branch adopts a U-Net architecture incorporating 2D convolutions and BEV-RWKV blocks (Fig. 4). Its encoder comprises an initial layer followed by four down-sampling stages, each integrated with a BEV-RWKV block. After processing the concatenated Semantic-BEV and Geometric-BEV features through the input layer and initial BEV-RWKV block, an ARF module [9] fuses these multi-scale representations to capture both semantic context and geometric structure. The decoder utilizes up-sampling operations and skip connections for spatial detail reconstruction, ultimately generating a 3D semantic occupancy grid , where denotes the class count.
@@ V CONCLUSIONS
In conclusion, OccRWKV, our novel network, successfully addresses the challenge of balancing performance and efficiency in 3D semantic occupancy prediction. It delivers state-of-the-art accuracy with a mIoU of 25.1 on the SemanticKITTI benchmark and maintains efficient real-time performance at 22.2 FPS. The network’s scalability makes it a robust solution for practical applications in robot navigation and autonomous driving. Field deployments confirm OccRWKV’s effectiveness in real-world settings, validating its suitability for future integration in complex environments.